OpenStudy (anonymous):

Find the exact value by using a half-angle identity. Sin(7pi/8)

OpenStudy (anonymous):

\(\frac{7\pi}{8}\) is half of \(\frac{7\pi}{4}\) so use the half angle formula for this one

OpenStudy (anonymous):

\[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\sin(x)}{2}}\] replace \(\sin(x)\) by \(\sin(\frac{7\pi}{4})\)

OpenStudy (anonymous):

@Hyun11 it does use latex, that is how you write here

OpenStudy (anonymous):

if you want to see the code, right click and choose "show math as" then "latex" you seem very angry relax

OpenStudy (anonymous):

so \[\sin (7\pi/4)/2)?\]

OpenStudy (anonymous):

no do you know what \(\sin(\frac{7\pi}{4})\) is ?

OpenStudy (anonymous):

The argument of trigonometric functions can be reduced modulo \[2\pi\]

OpenStudy (anonymous):

first find \(\sin(\frac{7\pi}{4})\) then replace that in the half angle formula here \[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\sin(x)}{2}}\]

OpenStudy (anonymous):

Write \[\frac{7\pi}{4}=\frac{8\pi}{4}-\frac{\pi}{4}=2\pi-\frac{\pi}{4}\]

OpenStudy (anonymous):

forget about "modulo \(2\pi\) " as \(\frac{7\pi}{4}<2\pi\)

OpenStudy (anonymous):

look at the unit circle, you will see that \[\sin(\frac{7\pi}{4})=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

Yeah so \[\sqrt{(1-7\pi/4)/2}\]

OpenStudy (anonymous):

then use the formula i wrote above

OpenStudy (anonymous):

no where you have \(\frac{7\pi}{4}\) you should have \(-\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

\[\sin(\frac{x}{2})=\pm\sqrt{\frac{1-\sin(x)}{2}}\] \[\sin(\frac{7\pi}{8})=\pm\sqrt{\frac{1-\sin(\frac{7\pi}{4})}{2}}\] =\[\pm\sqrt{\frac{1+\frac{\sqrt2}{2}}{2}}\]

OpenStudy (anonymous):

\[\sqrt{(1+2\pi/2})/2\]

OpenStudy (anonymous):

take a look at the answer i wrote above

OpenStudy (anonymous):

oh wait forget the pi >.<

OpenStudy (anonymous):

soo \[\sqrt{2-\sqrt{2}/2}\]?

OpenStudy (anonymous):

should be a plus sign inside the radical

OpenStudy (anonymous):

also there should be a minus sign outside of the whole thing, because \(\frac{7\pi}{8}\) is in quadrant 3 and therefore the sine is negative